one teacher has 4 times as many years experience as a second. six years from now, the first teacher will have twice as many years exprience as the second. how many years exprience does each teacher have now?
Answer by checkley77(12844)

Teacher 1=4x
Teacher 2=x
2(x+6)=4x+6
2x+12=4x+6
2x-4x=6-12
-2x=-6
x=-6/-2
x=3 years experience for Teacher 2.
4*3=12 years experience for teacher 1.
Proof:
2(3+6)=4*3+6
2*9=12+6
18=18